3. What would DuPonts net income for 2015 have been if it had been using FIFO to account for all of its inventories? In answCONSOLIDATED INCOME STATEMENTS Doillars in millions, except per share) For the year ended December 31 2015 2014 2013 Net saleInventory footnote question 3 December 31, 2015 2014 Finished products 3,779 $ 4,011 Semi-finished products 1,780 2,277 Raw mCan you explain why cash and equity change when the company switches from LIFO to FIFO, and why there is no change or up by 121.2 in net income and why does it different from the net income from Q3?
Answer #1

DuPont switched from LIFO to FIFO method to manage the inventories. To understand the the above changes first you need to understand what will happen when you switch from one method of inventory management to the other.

  1. Change in the closing inventory value Because while following LIFO method it is assumed that last inventories are sold first and inventory from the first batches remain in the closing inventory so the closing inventory was valued at the cost incurred to purchase/ produce the first batch of inventory.  

When you switch to FIFO method the reverse happens ( i.e) the inventories are sold in the order it is purchased/ produced. so the closing inventory will have goods which are purchased / produced lastly and it is valued at the cost incurred to purchase/ produce the last batch of inventory.  

So, Here it is given that Inventories went up by $202.

2. Similarly Cost of Goods sold will also change when you switch from one method to the other. As mentioned above under LIFO method the cost of goods would have been calculated by No. of units Sold * Cost of Last batch of inventories. But when you switch to FIFO method the the cost of goods should be calculated by No. of units Sold * Cost of first batch of inventories.

As we already know that the Closing inventories value went up by $202 which means the cost of goods sold would decrease by the same amount.

3. As we know that the cost of Goods sold is decreased by $202 which also means that Profit would have increased by same amount (i.e) $202.

4. Increase in Profit means increase in Taxes. So the we need to calculate the increase in Tax

Increase in Tax = 202 * 40% = $ 80.8

As the company is in 40% tax rate, we have calculated tax on increase in profit and the company would have paid tax more by $80.8 which is the reason for decrease in cash by $80.8

5. Increase in Profit before tax is $202. now lets find out increase in profit after tax.

Increase in Profit after Tax = 202 - 80.8 = $121.2

This Increase in Profit after tax of $121.2 means increase in Net income of the company. Net income after tax would have been added to Owner's equity which is the reason why equity is increased by $121.2.

2. Difference in Net Income from Q3:

In Q3 it is mentioned that Dupont follows FIFO for all inventories which means Change in Method of inventory has taken place for all years. But In Q4 changes is made only for 2015. As in Q4 changes is made only for 2015 we are changing the closing inventory value according to FIFO. If as mentioned in Q3 changes is made to all year then 2014 closing inventory value will increase by $240 ( inventory foot note ). which means increase in the value of opening inventory of 2015.

Increase in the value of Opening stock under FIFO method means increase in the cost of goods sold and increase in the value of closing stock under FIFO method means decrease in the cost of goods sold.

Cost of goods sold = Opening stock + Purchase/production - closing stock.

If FIFO method followed from 2014 means in 2015 opening stock value increases by $240 and Closing Stock value increases by $202. so, the Net effect the cost of goods sold increases by $38 ( 240 - 202). Increase in cost of goods sold means decrease in Net profit.

That is why net profit in Q3 is reduced by $38 less taxes.